Determine the area of the region bounded by x = y²-y-6 and x = 2y +4.

1
Expert's answer
2022-05-23T22:57:02-0400

y2-y-6=2y+4

y2-3y-10=0

y1=5

y2=-2


S="\smallint"-2 5 ((2y-4)-(y2-y-6)) dy = "\smallint"-25 3y-y2+2 dy = 1.5y2-y3/3+2y |-25 = 343/6

S = 343/6